1. Solid zinc was added to 1.0L of 0.100 M HCl and the following data were generated:

Time (seconds)

Mass of Zn (g)

0

0.016

4

0.014

8

0.012

12

0.010

16

0.008

a) Write a balanced chemical equation for this reaction.

b) Find the rate of reaction in moles of Zn consumed per time (units of moles/s)

c) Write the complete NET IONIC equation for this equation. (If you need help with writing a net ionic equation, see this video here)

d) What will happen as the [H+] decreases?

e) What will happen as the [Cl-] decreases?
Answer #1

a) Balanced chemical equation – Zn(s) +2H Cd (aq) → Znch (aq) + H₂ (g) (b) Initial moles of Zn = 0.016 - Cat t=o) 65 moles of

Rate of the reaction = 7.692 × 10 -6 mol/s